OpenStudy (anonymous):

What is the integral of e^-s^2ds from 0 to t

OpenStudy (anonymous):

(e^-s^2)ds

OpenStudy (anonymous):

The integral can not be written in terms of elementary functions.

OpenStudy (anonymous):

Oh this is for differential equations

OpenStudy (anonymous):

I need to find the derivative of y=(e^t^2 )(integral of e^-s^2ds from 0 to t)

OpenStudy (anonymous):

Then you need to use the fundamental theorem of calculus

OpenStudy (anonymous):

and the product rule.

OpenStudy (anonymous):

yeah but how about at the integral of e^-s^2ds from 0 to t part?

OpenStudy (anonymous):

The derivative of that is the integrand with t substituted for s

OpenStudy (anonymous):

derivative with respect to t, that is

OpenStudy (anonymous):

Okay so when I use product rule does the integral part when taking the derivative become e^t^2?

OpenStudy (anonymous):

In the other summand, you will not get around writing out the integral :D

OpenStudy (anonymous):

I think it'd be a lot more complicated than that?

OpenStudy (anonymous):

e^-t^2*

OpenStudy (anonymous):

it's 0 for some reason

OpenStudy (anonymous):

Nope it would be \[ e^{-t^{-2}} \]

OpenStudy (anonymous):

e^x is never 0. I don't know what you mean.

OpenStudy (anonymous):

is it becomes Im integrating in terms of t and not s

OpenStudy (anonymous):

Should I write out the whole derivative for you?

OpenStudy (anonymous):

sure I'm slightly confused

OpenStudy (anonymous):

Ok, first denote \[ f(t) = \int^t_0 e^{-1/s^2} ds \]

OpenStudy (anonymous):

I typed it in wolfram and got 0 and I remember my TA saying something like that

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then \[ f^\prime(t) = e^{-1/t^2} \]

OpenStudy (anonymous):

by the fundamental theorem of calculus

OpenStudy (anonymous):

Now set \[ g(t) = e^{t^2} f(t) \]

OpenStudy (anonymous):

Is this the function you wanted to know the derivative of ?

OpenStudy (anonymous):

Yeah but I don't see how it's e^-1/t^2?

OpenStudy (anonymous):

Well the fundamental theorem of calculus says: If \[ f(t) = \int^t_0 g(s) ds \] for some continous g, then f is differentiable and \[ f^\prime(t) = g(t) \]

OpenStudy (anonymous):

okay thanks for your help I shall ask again, he said this was a tricky one where most people would just sub in t for s

OpenStudy (anonymous):

just check out the fundamental theorem of calculus and use the product rule.. maybe the result is very simple, i didn't do the work

OpenStudy (anonymous):

Nope I just did the question and the answer for that has to be 0 it has something to do with it being t and the integral being in terms of s.